It has long been recognized that practical contexts for constraint satisfaction problems (CSP) often involve large relational databases (RDB). Early attempts to marry constraint solving systems and relational database systems include deductive and constraint databases that reuse important ideas from logic programming. These techniques required knowledge outside the scope of traditional database users. The recent proposal by Cadoli and Mancini, CONSQL, shows that a simple extension to SQL provides a viable basis for modeling CSP. This opens the possibility for transparently integrating CSP with databases using SQL the most widely known and popular database language. Such an extension brings the power of constraint problem solving to SQL knowledgeable users. Towards that end, the current research describes a case study in the engineering details of designing and implementing such a prototype.